Razor Inc. manufactures industrial components. One of its products used as a subcomponent in auto manufacturing is Fluoro2211. The selling price and cost per unit data for 9,110 units of Fluoro2211 are as follows. Per Unit Data Selling Price $ 370 Direct Materials 130 Direct Labor 26 Variable Manufacturing Overhead 23 Fixed Manufacturing Overhead 41 Variable Selling 14 Fixed Selling and Administrative 21 Total Costs 255 Operating Margin $ 115 During the next year, sales of Fluoro2211 are expected to be 10,110 units. All costs will remain the same except for fixed manufacturing overhead, which will increase by 20%, and direct materials, which will increase by 10%. The selling price per unit for next year will be $380. Based on these data, Razor Inc.'s total contribution margin for next year will be: (CMA adapted) Multiple Choice $1,561,140. $1,659,140. $1,651,140. $1,759,140.
Answer #1

Answer : $1,759,140

Next year Contribution MArgin :

Selling Price = 380

Direct MAterial for NExt year = 130*1.10 = 143

Total Varaible cost Per Unit = 143+26+23+14 = $206

Contibution MArgin per Unit = Selling Price - Varaible cost = 380-206= 174

Totalk contibution MArgin =10,110*174 = $1,759,140 (Answer)